OpenStudy (anonymous):

Solve 4x + 7 = 6x - 1

OpenStudy (uri):

Subract 7 both sides 4x + 7 = 6x - 1 -7 -7 ========== 4x=6x-8 4x-6x=8 -2x=8 x=8/-2 x=-4

OpenStudy (anonymous):

sorry i couldnt do eqn in original. its actually: \[4^{x+7} = 6^{x-1}\]

terenzreignz (terenzreignz):

Get the log of both sides. It's the only way...

OpenStudy (anonymous):

so log(x+7) 4 = log(x-1) 6?

terenzreignz (terenzreignz):

Let's use natural logarithm. \[\Large \ln(4^{x+7}) = \ln (6^{x-1})\]

terenzreignz (terenzreignz):

Hmm? Use this property \[\Large \ln b^p = p\ln (b)\]and the fact that ln(4) and ln(6) are just constants.

OpenStudy (anonymous):

how do i enter that into my calculator?

terenzreignz (terenzreignz):

Not yet. You still have to solve.

terenzreignz (terenzreignz):

By isolating x.

OpenStudy (anonymous):

so get x on one side?

terenzreignz (terenzreignz):

Yup. But you still have to deal with these:\[\Large \ln(4^{x+7}) = \ln (6^{x-1})\] and use that property I posted.

OpenStudy (anonymous):

im not sure how

terenzreignz (terenzreignz):

The exponent just becomes a factor... \[\Huge \ln b^{\color{red}p}= \color{red}p\ln(b)\]

OpenStudy (anonymous):

so it would become ln(4^[x+7]) =(x-1)ln(6)?

terenzreignz (terenzreignz):

Yes... partly, but you should do the same to the left side :P

OpenStudy (anonymous):

oh i thought it was like how the equation was set up, sorry about that so (x+7)ln(4)=(x-1)ln(6)?

terenzreignz (terenzreignz):

Yup. And after a bit of algebraic manipulation, you get \[\Large \frac{x+7}{x-1}=\frac{\ln 6}{\ln 4}\] \[\Large \frac{x+7}{x-1}=\color{blue}{\log_46}\] and that part in blue is just a constant you can enter into your calculator (if not, then just divide ln 6 by ln 4)

OpenStudy (anonymous):

so x+7/x-1 =~ 1.8

terenzreignz (terenzreignz):

yes, and just solve normally.

terenzreignz (terenzreignz):

I don't know what log(base4) of 6 is so I'll just leave it as k for now. You have the calculator, you know what it is \[\Large x+7 = \color{blue}k(x-1)\] And this is just your basic linear equation :P Solve as you see fit XD

OpenStudy (anonymous):

x+7~1.8(x-1) x+7~1.8x-1.8 -.8x~5.2 x~-6.5 is what i got?

terenzreignz (terenzreignz):

Nope. You subtracted 7 from 1.8x They're not similar terms. Never forget your basics.